OpenStudy (anonymous):

Solve for x: |3x + 2| = 14

OpenStudy (anonymous):

x=4 or -5.3 im not sure witch one

OpenStudy (anonymous):

for positive charge 3x+2=14 =>3x=14-2 =>3x=12 =>x=4 for negative charge -(3x+2)=14 =>-3x-2=14 =>-3x=14+2 =>-3x=16 =>x=-16/3

OpenStudy (anonymous):

*which